two men start togather to walk a certain distance. one at 4km/hr and the other at 4/3 km/h . the former arrives 30 minutes before. what is the total distance in kilometers?
Answer by mananth(16949)

two men start togather to walk a certain distance. one at 4km/hr and the other at 4/3 km/h . the former arrives 30 minutes before. what is the total distance in kilometers?
...
let distance be x
..
Man A time = x/4 hours
Man B time = x/(4/3) hours
Time of B - Time of A = 1/2
..
3x/4 - x/4 = 1/2
2x/4 = 1/2
4x = 4
x= 1 km ...... Distance